OpenStudy (anonymous):

Hi all! How do I get the x-intercept of -x^3 +3x^2 +2 ? I understand you set it equal to zero, but after that I'm stuck :/

OpenStudy (anonymous):

so basically u got stuck while solving the cubic equation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So I have -x^3 +3x^2 =-2

OpenStudy (anonymous):

then factor?

OpenStudy (anonymous):

No, because it's not equal to 0

OpenStudy (anonymous):

hmm

ganeshie8 (ganeshie8):

-x^3 + 3x^2 = -2 \(x^3 - 3x^2 -2 = 0\) eliminate x^2 term, put y = x-1 => \(x = y+1\)

ganeshie8 (ganeshie8):

this is going to be a long process, u may give up if u dont wanto deal wid all the mess + complex numbers that wil come on the way.. :)

ganeshie8 (ganeshie8):

u can simply approximate the root wid newton if u want

OpenStudy (anonymous):

Hmm, I don't know this method

OpenStudy (anonymous):

how do I do it?

ganeshie8 (ganeshie8):

pick one below :- 1) newton approximation (calculus) 2) algebraic solution

OpenStudy (anonymous):

1

ganeshie8 (ganeshie8):

very wise :)

OpenStudy (anonymous):

How do I start?

OpenStudy (anonymous):

like this? http://www.youtube.com/watch?v=1uN8cBGVpfs

ganeshie8 (ganeshie8):

yes, lets start with x = 3 maybe

OpenStudy (anonymous):

yeah, I was thinking between 3.1 and 3.2, is that good?

ganeshie8 (ganeshie8):

\(f(x) : x^3 - 3x^2 -2 \)

OpenStudy (anonymous):

ah ok

ganeshie8 (ganeshie8):

\(\large x_{n+1} = x_n - \frac{f(x_n)}{f'(x_n)}\)

ganeshie8 (ganeshie8):

start wid \(x_0 = 3\)

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

its a recursive formula, the approximation gets better as the iterations increase

OpenStudy (anonymous):

mhm

OpenStudy (anonymous):

is that the algebraic solution?

OpenStudy (anonymous):

i used wolfram coz it was beyond my level .!

ganeshie8 (ganeshie8):

yep, u dont want to do that.. not in ur scary night mares :|

OpenStudy (anonymous):

XD it's scaaary

OpenStudy (anonymous):

so, the newton method will yield the same/a similar result?

ganeshie8 (ganeshie8):

lets see :) \(\large x_{n+1} = x_n - \frac{f(x_n)}{f'(x_n)}\) \(f(x) : x^3 - 3x^2 -2 \) \(f'(x) : 3x^2 - 6x\) Iteration 1 : \(\large x_1 = 3 - \frac{f(3)}{f'(3)}\)

ganeshie8 (ganeshie8):

thats the approximation after first iteration

OpenStudy (anonymous):

I don't understand this part of the notation which says frac

ganeshie8 (ganeshie8):

oh you're not seeing latex, there is a problem with latex in this site for some users

OpenStudy (anonymous):

ah, what must I do?

ganeshie8 (ganeshie8):

come here http://www.twiddla.com/1462458

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thanks again everyone for your help!

ganeshie8 (ganeshie8):

u wlc :)